OpenStudy (anonymous):

Derive ln(tanx) Howdo I do this? I tried the chain ruled but did not get the right answer

OpenStudy (amistre64):

can you show what you did?

OpenStudy (amistre64):

recall that ln(u) derives to u'/u

OpenStudy (jdoe0001):

\(\bf \cfrac{d}{dx}[{\color{blue}{ ln[{\color{brown}{ tan(x)}}]}}]\implies {\color{blue}{ \cfrac{dy}{dx}}}\cdot {\color{brown}{ \cfrac{dy}{dx}}}\)

OpenStudy (anonymous):

this is what I did. u=tanx ln(u)*u' so I got sec^2x/tan(x)

OpenStudy (amistre64):

sec^2/tan is good ..... it can be reworked of course, but its fine

OpenStudy (jdoe0001):

yeap.... maybe what you're looking at is some simplified version of that

OpenStudy (amistre64):

\[\frac1{c^2}\frac{c}{s}=\frac{1}{cs}\]

OpenStudy (amistre64):

sec cos maybe?

OpenStudy (amistre64):

lol, csc not cos inmy mind it was right but my fingers hate me

OpenStudy (jdoe0001):

yeap.... could be rewritten as csc(x)sec(x)

OpenStudy (aum):

What is the "right answer" that you are referring to?

OpenStudy (amistre64):

hope its not x^2 :)

OpenStudy (anonymous):

in my book it says the answer is 2csc2x

OpenStudy (amistre64):

well sin(x) cos(x) = 2sin(2x) but then thats csc(2x)/2

OpenStudy (aum):

\[ \frac{\sec^2(x)}{\tan(x)} = \frac{1}{\cos^2(x)} * \frac{\cos(x)}{\sin(x)} = \frac{1}{\sin(x)\cos(x)} = \frac{2}{2\sin(x)\cos(x)} = \frac{2}{\sin(2x)} = 2\csc(2x) \]

OpenStudy (amistre64):

so what your saying is trig is a horrible subject and should simply be eliminated from all cirrculum .... i agree :)

OpenStudy (anonymous):

help me I am confused! how do I get that

OpenStudy (amistre64):

you have to remember your trig indentities